### v1

# Assessing Physics Students’ Scientific Argumentation using Natural Language Processing

Scientific argumentation is an important science and engineering practice and a necessary 21st Century workforce skill. Due to the nature of large enrollment classes, it is difficult to individually assess students and provide feedback on their argumentation. The recent developments in Natural Language Processing (NLP) and Machine Learning (ML) may provide a solution. In this study we investigate methods using NLP and ML to assess and understand students argumentation. Specifically, we investigate the use of topic modeling to analyze student essays of argumentation after solving a problem in the recitation section of an introductory calculus-based physics course four semesters. We report on the emergent themes present in each semester.
Assessing Physics Students’ Scientific Argumentation using Natural Language Processing
Winter Allen Department of Physics and Astronomy, Purdue University, 525 Northwestern Ave, West Lafayette, IN 47907, U.S.A.    Carina M. Rebello Dept. of Physics, Toronto Metropolitan University, 350 Victoria Street, Toronto, ON M5B 2K3, Canada    N. Sanjay Rebello Dept. of Physics and Astronomy / Dept. of Curriculum & Instruction, Purdue University, West Lafayette, IN 47907, U.S.A.
1 Introduction
For most students, their first official introduction to physics is in a high school classroom. This initial exposure often leaves a lasting impression, shaping how they view the subject for the rest of their lives. Many people walk away viewing physics as hard and not relevant to their day-to-day lives. However, learning physics encompasses far more than solving equations; it is about fostering a way of thinking that helps us understand the natural world. Oftentimes the disconnect between these aspects can be detrimental to student learning of physics. A core practice we want students to develop is the ability to solve problems; yet, the focus on plugging equations often overshadows this broader educational goal.
Student problem solving is a relevant and rich topic in Physics Education Research (PER) and has been for many years [1, 2, 3]. Students often prioritize memorizing final answers over developing a deeper understanding of the problem-solving process [2, 4]. To foster growth in problem-solving, it is crucial not only to understand why students solve problems the way they do but also to help them reflect on their own problem-solving strategies, allowing them to develop as both learners and future scientists.
Several research-based strategies have attempted to address the issues students have with problem solving [1]. We aim to address these fundamental issues through scientific argumentation. This idea is deeply rooted in philosophy [5] and has evolved significantly through educational research [5, 6]. Scientific argumentation is a proven strategy to help improve critical thinking that provides a schema for justifying the relevance of the retrieved knowledge in problem solving. To construct an argument students must justify their methods and decisions as they solved a problem, go through every step they took up to their solution, and provide evidence and reasoning for their process. In the context of problem-solving in physics, scientific argumentation involves not only an explanation of conceptual knowledge and methods but the ability to justify reasoning with empirical evidence and logical consistency. Within PER, scientific argumentation has been shown to enhance students’ ability to link theoretical knowledge with practical problem-solving skills [7]. This process encourages students to think critically about the methods they use and the evidence they gather, promoting skills that are essential for expert-like problem solving. Peer argumentation in physics classrooms also fosters collaborative learning, where students refine their ideas through group discussions and critique, further advancing their conceptual understanding and reasoning abilities [8]. The process of reflecting and evaluating one’s solution that the iterative nature of argumentation provides aligns with the goals of PER in promoting both content mastery and the development of scientific critical thinking.
Scientific argumentation can be studied in both oral and written modalities. In this work, we focus on investigating scientific argumentation in students’ written essays in which they describe their strategies for solving problems. One of the major challenges that we face is systematically teaching and assessing student written essays in large enrollment classes due to the prohibitive time it takes for instructors or teaching assistants to read students’ written work and gauge their argumentation quality. Recent developments in Large Language Models (LLMs), Natural Language Processing (NLP), and Machine Learning (ML) may afford us the opportunity to address these challenges. ML is a subset of Artificial Intelligence (AI) that is the development of algorithms to detect patterns in datasets. NLP and LLM’s are subsets within ML that focus on and aid in the understanding of human language. Recently, there have been many studies on using LLMs, NLP, and ML in assessing student writing. [9, 10, 11, 12, 13]. Researchers have used these tools, such as unsupervised NLP, to assess student strategy essays [13] and explore the viability of utilizing LLMs in physics education [12, 11, 9, 10]. NLP and ML provide unique tools for analyzing large amounts of student text, allowing for the identification of patterns while still enabling qualitative, in-depth studies.
To approach assessing students’ argumentation with ML in the context of problem-solving we have designed a study that introduces scientific argumentation through a series of scaffolds in the recitation portion of the course. We offer increasing levels of scaffolds through a four semester study. We then utilize unsupervised machine learning techniques to assess how student arguments change through the semesters.
2 Research Questions
The specific research questions (RQs) are:
RQ1: By introducing more levels of scaffolding throughout the semester, to what extent will students develop more complete, thoughtful arguments by the last module?
RQ2: To what extent can we use machine learning methods to assess student argumentation in the context of physics problem-solving?
3 Background
3.1 Scientific Argumentation
We live in an age where misinformation is spread with a click of a button. We are constantly bombarded with dissenting information requiring strong critical thinking skills to wade through to the truth. Students often view science as another subject to memorize and regurgitate facts, rather than as a dynamic, evolving body of knowledge. This fundamental issue can be addressed by encouraging students to think critically about their work and engage in Scientific Argumentation (SA) in a classroom. It has been shown that by constructing explanations students may change their view of the nature of science [14]. As scientists, we know it is important to critically reflect on our work. Whether it be in research or solving a simple freshman level problem. We not only need to understand what decisions we made and why we made them in the problem-solving process, but we need to be able to present a well supported argument in support of our process. Research suggests that students tend to struggle with the idea of developing scientific arguments[15, 16], especially with finding appropriate evidence and constructing their reasoning [17, 18] and distinguishing between various elements of an argument [19, 20].
Despite these difficulties, truly engaging in argumentation has shown to be beneficial to students. Not only does it aid in their own understanding of the problem, but it helps them learn to communicate and support their own findings [21]. To aid students in constructing arguments: argumentation scaffolds can elicit students’ participation in scientific argumentation [22] and a conducive learning environment can support students to solve problems, compare solutions, consider alternatives, and justify choices [17], [20]. Appropriate scaffolds include justification prompts [23] and question prompts [24]in instructional materials that help students articulate the rationale for their problem-solving steps and urge them to reason using evidence and justifications [25, 6] based on underlying principles [26]. However, most undergraduate physics courses do not facilitate scientific argumentation. Curricula that facilitate more expert-like problem solving can positively influence students’ epistemic beliefs and expectations around problem solving [27]. In more recent work, Rebello et al. [28, 7] found positive effects of using scientific argumentation in physics courses for future elementary teachers as well as future engineers.
In recent years, scientific argumentation has been studied in various subjects: biology [29] [30], chemistry [31], and physics [32]. There are multiple ways to assess scientific argumentation [33] that rely on context in which the argument is presented, nature of the task being performed, and the specific learning goals [34]. For example, scientific argumentation in inquiry-based tasks may be assessed differently than in conceptual problem-solving activities. Moreover, these assessments must account for the complexity of scientific reasoning, the use of evidence, and how students articulate and justify their claims [35]. As a result, educators and researchers have developed diverse methodologies, such as Toulmin’s argumentation model [36, 5], which emphasize the structure of arguments, as well as rubrics that measure the quality of evidence and reasoning in students’ responses [37].
Toulmin, a British philosopher, proposed breaking down arguments into six components: claim, grounds, warrant, qualifier, rebuttal, and backing.[5] A claim is the base purpose of an argument. The grounds are the evidence of the argument that support the claim. The warrant links the grounds to the claims. This can be explicitly stated or implied. The qualifier is used in wording the claim, while the rebuttal and backing are implied. The rebuttal acknowledges things that may contradict the claim. The backing establishes the relevance of the warrant. [36], [38] Toulmin’s model is based on arguments present primarily in law; however, this model has been particularly useful in examining how students construct arguments and justify their claims.
McNeill and Krajcik [39] further adapted Toulmin’s model for use in science education by simplifying it into the Claim-Evidence-Reasoning (CER) framework. In their framework, the claim is an assertion or conclusion about a phenomenon, the evidence consists of scientific data supporting the claim, and the reasoning explains the relevance of the evidence. CER has become a popular framework in K-12 education, where students are encouraged to construct arguments using data to support their claims [40, 41]. Given the effectiveness of CER in K-12, there is a strong rationale for exploring its adaptation in undergraduate physics education, where developing students’ ability to argue scientifically can enhance their problem-solving and critical thinking skills.
Scientific argumentation has been a rich topic to study in science education in K-12 classrooms in recent years. Erduran and Park [42] performed a search for manuscripts focusing on argumentation in secondary PER between 2003 and April 2022. During this period, they identified 13 published studies that explored scientific argumentation in secondary physics classrooms. These studies have provided critical insights into how students develop argumentation skills at the K-12 level, emphasizing the importance of scaffolding and structured support for helping students engage with complex scientific reasoning. However, despite the growing interest in this area, the relatively small number of studies over two decades suggests that research on argumentation, particularly in physics, is still in its early stages. Furthermore, these findings highlight the need for continued efforts to expand scientific argumentation research, particularly given the increasing emphasis on argumentation as a critical component of scientific literacy in modern curricula [37].
Even less work has been done in undergraduate physics classrooms. While Erduran and Park [42] found 13 manuscripts on argumentation in secondary PER from 2003 to 2022, only nine manuscripts were published during the same period focusing on tertiary PER, highlighting the gap in research at the undergraduate level. In these nine manuscripts, there was a range of focus. The lack of study in this context does raise concerns as argumentation plays a critical role in developing advanced scientific reasoning and expert-like critical thinking skills, which are essential for physics students at the tertiary level as they transition into more complex problem-solving tasks that will aid them in the 21st century job market.
In these nine manuscripts, a range of focus areas emerged. Some research concentrated on the epistemic tools learners use to construct scientific arguments [43], investigating how students draw upon their knowledge, reasoning, and evidence to build coherent, well-supported arguments. Other studies examined the relationship between student content knowledge and their performance on argumentation tasks [44], showing that students’ ability to engage in scientific argumentation often hinges on their depth of understanding of the underlying physical principles. Additionally, some research has explored the ways in which students use arguments across different disciplinary contexts, such as mathematics and physics, highlighting the challenges students face when applying argumentation strategies in varied subject areas [45].
Despite these valuable insights, the limited volume of research on scientific argumentation in undergraduate physics classrooms points to the need for continued research in the topic. There is a lack of comprehensive studies that examine how scientific argumentation can be systematically integrated into traditional physics curricula. Most undergraduate physics courses tend to emphasize problem-solving over argumentation, leaving little room for students to engage in the kind of reflective, evidence-based reasoning that scientific argumentation requires. We aim to contribute by exploring how scientific argumentation can be effectively integrated into undergraduate physics recitations to improve students’ arguments.
3.2 Machine Learning
Machine learning is a branch of artificial intelligence that focuses on the development of statistical algorithms to understand patterns in unseen data. Two general sets of machine learning are unsupervised and supervised. Supervised machine learning relies on known, labeled data to train an algorithm to predict patterns in unseen data. Unsupervised machine learning focuses on discovering hidden patterns within datasets without explicit human guidance (training). Some popular examples of machine learning are clustering and topic modeling. In this paper, we chose to focus on unsupervised machine learning to analyze our data, due to to the large amount of student responses it would not be feasible to qualitatively extract emergent themes; nor would it be possible to label enough data to approach this project in a supervised or even semi-supervised way. We have worked with students for multiple years on argumentation, specifically on the same set of recitation problems, so we are confident in what we expect to see for the average student argument. Therefore, we feel confident in approaching this dataset with unsupervised techniques.
(i) Topic Modeling
Topic modeling is a technique that falls within the field of unsupervised machine learning. It is used to uncover hidden thematic structures in large collections of text. Rather than relying on labeled data, topic modeling algorithms infer topics based on patterns of word co-occurrence across documents.
One popular method of topic modeling is Non-Negative Matrix Factorization (NMF) [46], which is what we chose to use in this study. NMF is rooted in linear algebra and works by approximating an original document-term matrix as the product of two lower-dimensional matrices. The key constraint in NMF is that all values must be non-negative, which tends to produce more interpretable results—especially useful when analyzing human language, where negative values lack intuitive meaning.
Basically, suppose we have a document-term matrix $V$ of size $m\times n$, where $m$ is the number of documents and $n$ is the number of unique terms. NMF factorizes $V$ into two non-negative matrices: $W$ and $H$. $V$ represent the term document matrix, $W$ is an $m\times k$ matrix (where $k$ is the number of topics). Each row in $W$ represents a document’s distribution over topics. $H$ is a $k\times n$ matrix, where each row corresponds to a topic’s distribution over the terms. The idea is to find non-negative matrices W and H such that their product closely approximates the original matrix $V$. The rows of matrix $W$ give the document-to-topic weights, i.e., how much each document is associated with each topic. The rows of matrix $H$ give the topic-to-term weights, i.e., how much each term contributes to a given topic.
4 Methodology
4.1 Dataset
This study was implemented in a first-semester calculus-based physics class for future engineers, at a large U.S. Midwestern land grant University. The annual enrollment at the time of the study is approximately 2500 students (1100 in fall and 1400 in spring). The course is built around three key principles: momentum, energy, and angular momentum, and it follows Chabay and Sherwood’s Matter & Interactions [47]. The course has three components: lectures (two 50-minute-long session), lab (one 110-minute-long session, and recitations (one 50-minute-long session). The recitations are the context of this study.
The recitations are led by one graduate teaching assistant (GTA) with the aid of an undergraduate teaching assistants (UTA). During each Recitation session, the GTA’s spend a short time at the beginning of the recitation introducing the problem and relevant information. They start off each section by going over a provided Powerpoint presentation introducing the recitation problem. Each presentation takes around 5 to 10 minutes. In earlier weeks of the semesters, students receive more help on starting the problem, such as going over system and surroundings as a class. As the semesters progress students receive more general help on understanding the problem and less help on starting the practical aspects of it. Then students are expected to work together in groups of 3-5 at their table to solve the problem. They often use a whiteboard to create a collaborative workspace to discuss and share their work. As students work together, the GTA and UTA are available and walking around the room to help groups when they have questions. Even though they work in groups, every student is expected to provide their answers to questions in an individual Jupyter Notebook file (.ipynb extension). They edit this file on Google Colab in class. Students discuss the questions collaboratively and enter their answers to these questions and input an image of their written work.
After they are finished, they are expected to submit the file and a PDF of the file into Brightspace. The notebook file is uploaded so we have access to easily extract student responses. A PDF is uploaded for grading purposes. Students are graded as a group by GTAs on a rubric provided. GTAs choose one student randomly from each group, grade them for correctness, and assign everyone in the group this grade. Students are not expected to spend much, if any, time on the recitation outside of the 50-minute section, so there should be very little individual work on the recitation other than expressing their answers in their own words. Due to the nature of the grading, students are encouraged to ask the TAs questions and make sure everyone in the group is on the same page before they leave the Recitations session.
Figure 1: Recitation Problem for Student Argumentation
4.2 Scaffolding
Throughout the last two years, students were tasked with writing argumentation essays at the end of most recitations. Table 1 outlines how the scaffolds progressed each semester. The example question in the second row of each column represents the question students were asked at the end of the semester after receiving their full set of scaffolds. The dataset we report on is in response to question in the table.
The study started in Fall 2022 when students received no instruction on argumentation, except of the prompt shown in Table 1. In Spring 2023, we implemented argumentation prompts based upon McNeil and Krajcik’s model within the recitation. We gave students the definitions of claim, evidence, and reasoning (CER) in pieces as the semester progressed, while prompting them to list their own CER in parts before writing a full argument of their own. Finally in Fall 2023 and Spring 2024 students were given similar scaffolds to that of Spring 2023. However, instead of immediately having students list their own CER, we gave students statements and had students answer which statements were CER. Then in the middle portion of the semester students listed their own CER. Finally they constructed their own arguments at the end of the semester.
Table 1: Argumentation Scaffolding
Scaffolding Fall 2022 Spring 2023 Fall 2023 / Spring 2024
Module 1 Students were asked to construct arguments Students were given definitions of Claim, Evidence, and Reasoning (CER) and asked to list their CER, separately. Students were given definitions of CER and prompts to identify CER statements.
Module 2 Students were asked to construct arguments Students were given definitions of CER and asked to construct arguments. Students were given definitions of CER and asked to list their CER, separately.
Module 3 Students were asked to construct arguments Students were given definitions of CER and asked to construct arguments. Students were given definitions of CER and asked to construct arguments.
Study Question In words, construct an argument to explain and justify your solution. Justify the various decisions you took while constructing your solution. In your argument, incorporate why the chosen principle(s)/concept(s) and assumption(s)/approximation(s) are relevant to your proposed solution. In words, construct an argument to explain, elaborate and justify your solution. Your argument should be in a paragraph and contain the CLAIMS, EVIDENCE, and REASONING that support your solution. In words, construct an argument to explain, elaborate and justify your solution. Your argument should be in a paragraph and contain the CLAIMS, EVIDENCE, and REASONING that support your solution.
4.3 Topic Modeling
We analyzed open-ended student scientific arguments from each of the four semesters separately. Although, students were given instruction to use complete only words and complete sentences to construct their essays, there still required cleaning of the text. We began by employing a comprehensive text pre-processing pipeline to clean student argumentation essays. This included punctuation and number removal, conversion to lowercase, and stopword (e.g. ’the’, ’and’ , ’as’, etc.) filtering using NLTK [48]. Additionally, we implemented automated spell checking using the pyspellchecker library to correct common typos. Responses with fewer than 10 words were filtered out to reduce noise and trivial responses.
We applied non-negative matrix factorization (NMF) to extract latent topics from the student essays. First, the pre-processed text was vectorized using the TF-IDF method, capturing term importance across the corpus. We then applied the NMF algorithm [46] with ten components (topics). Each component represents a distribution over words, and each essay is represented as a mixture of topics.
To identify the optimal number of topics for NMF, we evaluated both the reconstruction error [46] and coherence scores across topic counts ranging from 2 to 30. The reconstruction error, which measures the discrepancy between the original TF-IDF matrix and its reconstruction from the NMF model, was calculated for each topic configuration. This metric is used to assess how well the NMF model approximates the original data. Coherence scores provide an indication of how interpretable the topics are. We used the Coherence Model [49] from the gensim library to compute coherence scores for the topics generated by the NMF model. As expected, reconstruction error decreased monotonically with additional topics, Figure 2, reflecting better matrix approximation. However, coherence scores, which measure the semantic interpretability of topics, fluctuated and did not exhibit a clear maximum . Based on a balance between interpretability and parsimony, and in line with prior work using similar data, we selected 10 topics as a reasonable compromise for downstream analysis.
(a) Fall 2022
(b) Spring 2023
(c) Fall 2023
(d) Spring 2024
Figure 2: NMF Reconstruction Error vs Number of Topics
To interpret the topics, we examined the top-weighted words associated with each component. Each essay was assigned to its most dominant topic using the argmax of the topic mixture vector. We visualized topic distribution across the corpus using a histogram, highlighting the number of student essays per topic. To compare the semesters, we analyzed the representative words of each topic, the number of essays in each topic, and performed a concentration analysis [50] of each semester. Bao and Redish [50] introduced this method to measure how students’ responses on multiple-choice questions are distributed. We have applied this method to measure how our student essays are distributed over different topics.
$C=\frac{\sqrt{m}}{\sqrt{m}-1}\times\frac{\sum_{i=1}^{m}n_{i}^{2}}{N}-\frac{1}{\sqrt{m}}$
Where, $m$ represents the number of topics, $n_{i}$ represents the total number of students who selected topic $i$, and $N$ represents the total number of students in the semester. A high concentration factor $C\approx 1$ means most essays are concentrated in one or a few topics, where as $C\approx 0$ suggests an even spread across all topics.
5 Results and Discussion
We report on the top five words of each topic and the essay distribution over topics of each semester. The resulting top five words were calculated to assess what each topic represented (2,3,4,5), and the number of essays in each topic are represented in the histograms (3(a),3(b),3(c), 3(d)) shown below.
Topics 1 2 3 4 5 Counts
Topic 0 energy kinetic change work equal 96
Topic 1 acceleration torque net angular force 73
Topic 2 system extended point particle part 80
Topic 3 solution correct made us sure 56
Topic 4 force applied cylinder work distance 90
Topic 5 mass center velocity force solve 64
Topic 6 used solved part energy principle 60
Topic 7 air resistance problem assumed cylinder 66
Topic 8 equation using part solve theorem 105
Topic 9 find velocity use final principle 51
Table 2: Top 5 words for each topic in Fall 2022 using NMF
Topics 1 2 3 4 5 Counts
Topic 0 energy kinetic translational rotational potential 71
Topic 1 cylinder incline slipping without rolling 84
Topic 2 used part principle find momentum 88
Topic 3 force friction applied distance gravity 98
Topic 4 system point particle extended wheel 66
Topic 5 mass center velocity find using 59
Topic 6 work change done energy equal 58
Topic 7 equation acceleration answer get vim 86
Topic 8 problem evidence claim earth surroundings 90
Topic 9 use principle solve energy need 65
Table 3: Top 5 words for each topic in Spring 2023 using NMF
Topics 1 2 3 4 5 Counts
Topic 0 system extended particle point part 68
Topic 1 energy kinetic work done equal 170
Topic 2 acceleration momentum angular linear torque 80
Topic 3 force applied distance twice friction 83
Topic 4 use energy principle problem must 87
Topic 5 friction work contact point move 72
Topic 6 object pulling problem would large 45
Topic 7 disk upwards slipping force counteracts 18
Topic 8 cylinder ramp slipping belt earth 96
Topic 9 used find velocity equation mass 113
Table 4: Top 5 words for each topic in Fall 2023 using NMF
Topics 1 2 3 4 5 Counts
Topic 0 system particle point extended friction 87
Topic 1 cylinder slipping without incline force 137
Topic 2 energy kinetic work change done 178
Topic 3 momentum angular principle find use 73
Topic 4 ramp surroundings earth belt system 94
Topic 5 force applied friction equal times 116
Topic 6 velocity mass center find using 78
Topic 7 work displacement ball slip slipping 47
Topic 8 cylinders speed workenergy theorem motion 39
Topic 9 solve part principle energy used 124
Table 5: Top 5 words for each topic in Spring 2024 using NMF
In fall 2022, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "force" and "part" were repeated three times while "energy", "work", "cylinder", "velocity", "solve", and "principle" were repeated twice. Across the topics, there were three relative peaks in Topics 0, 4, and 8 with the largest peak at Topic 8. Topic 0 focuses on concepts from the energy principle such as "energy", "kinetic, and "work. Topic 4 focuses on superficial features such as cylinder", "applied", "distance" and ideas of work such as "work" and "force". Topic 8 focuses on apparent mathematical steps to solve the problem which is apparent with the words "equation", "using", "part", "solve", and "theorem". Even though there are three peaks, the topics seem to be evenly distributed, meaning there is a broad range of topics that students focused on equally.
In Spring 2023, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "energy" was repeated 3 times while "find" and "principle" were repeated twice. Across the topics, there were two relative peaks in Topics 3 and 8 with the largest peak at Topic 3. Topic 3 focuses primarily on the forces such as "friction", "gravity", and literally "force". "Applied" likely is in reference to applied force. Topic 8 focuses not on physics concepts but argumentation terms such as "evidence", "claim, and generic "problem". Even though there are two peaks, the topics seem to be evenly distributed, meaning there is a broad range of topics students equally focused on.
In fall 2023, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "energy", "point", "work, "force", "friction", "problem" and "slipping" were repeated twice. Across the topics, there was one relative peak Topic 1. Across the topics, there was one relative peak in Topic 1, which included 170 essays. Topic 9 had the second most number of essays at 113. Topic 1 focuses on words we would expect from the energy principle, such as "energy", "kinetic", and "work". In this semester, there seems to be an unequal distribution of topics with essays skewing towards Topic 1.
In spring 2024, we found the top five words for each topic and the number of essays allocated to each topic. Within the topics there were some repetition of words between topics: "energy", "system, "friction", "slipping", "work", "force", "principle", and "find" were repeated twice. One odd outlier that appeared was a conjunction of two words "workenergy". Across the topics, there was one relative peak Topic 2, which included 178 essays. Topic 1 had the second most number of essays at 137. Much like in fall 2023, topic 2 focuses on words we would expect from the energy principle, such as "energy", "kinetic", and "work". In this semester, there seems to be an unequal distribution of topics with essays skewing towards Topic 2.
(a) Fall 2022
(b) Spring 2023
(c) Fall 2023
(d) Spring 2024
Figure 3: Number of Essays in Each NMF Topic across Four Semesters
After investigating the top words and distribution, we found the first two semesters, which had either no or less scaffolding, seemed to be evenly distributed over topics. Which means as a whole we had students focusing on important and correct concepts, such as energy, but equally focusing on trivially concepts, such as mathematical steps. In the last two semesters, in which students received the most scaffolding, we saw unequal distributions, with peaks in topics that focused on energy ideas. This indicates that students in these semesters tended to focus more on physics oriented ideas.
Figure 4: Concentration Factors over Four Semesters
To further investigate the distribution across topics, we performed a concentration analysis and report on the coherence factor in each semester in Figure 4. We see an increase in concentration from Fall 2022 to Spring 2024. The Fall 2023 and Spring 2024 semesters have higher concentrations of essays than Fall 2022 and Spring 2023 semesters. It is interesting to note we may begin to see a pattern emerging between Fall and Spring semesters. The concentration of Spring 2023 is less than that of Fall 2022, and the concentration of Spring 2024 is less than that of Fall 2023. More investigation needs to be done to determine if this pattern holds for more semesters; however, there could be interesting implications to this finding. When students have misconceptions it tends to result in more diverse incorrect answers, while correct student answer tend to reflect more uniform understanding [51]. Therefore, in Spring semesters students may have more misconceptions than students in fall semesters, since we see less coherence.
6 Conclusions
To address our first research question: by introducing more levels of scaffolding throughout the semesters, we found that students constructed arguments focused more on physics concepts opposed to mathematical steps. To address our second research question: we were able to use unsupervised NLP methods to assess student argumentation in the context of physics problem-solving.
By investigating the top words of the topics and the counts of essays in each topic, we were able to see a change in student focus across four semesters. In the semesters, with less or no scaffolding students seemed to focus equally across topics, with some topics focusing on superficial features, mathematical concepts, or physics concepts. In the latest two semesters, we saw more essays focus on energy principle concepts than any other topic in the respective semesters. In these semesters, students received the same levels of scaffolding, which was more than the previous two. This supports that we were able to improve student argumentation and track the progress over the course of two years. In addition, we measured the coherence factors and see that later semesters have more coherence that that of Fall 2022.
6.1 Implications and Future Work
The results of this study show promise for exploring various unsupervised machine learning approaches to compare students across different semesters. We intend to expand this work to track student scientific arguments throughout a single semester to determine if we can capture their argumentation progress using unsupervised machine learning techniques.
7 Acknowledgments
This research is supported in part by U.S. National Science Foundation grants 2111138 and 2300645 Any opinions, results, and findings expressed here are those of the authors and not of the Foundation.